From nch@roe.ac.uk Fri Oct 9 17:50:38 2009 Date: Fri, 9 Oct 2009 16:27:05 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 09/10/09 Minutes of WFAU VDFS Science Archive meeting: October 9th 2009 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: RSC, PMW, ETWS, RPB, NCH, MAR, NJC Apologies: JPE, JMS, AL, RGM, MSH DoNM: 10am, Friday October 16th 2009 in the VISTA Hut Actions discharged this week: ----------------------------- ACTION: RPB to investigate share /mnt on ramses2 Discharged; RSC is testing ... ACTION: NCH & PMW to meet Monday 2pm to flesh out the Q4 plan Discharged; see below Actions partly discharged but continuing: ----------------------------------------- ACTION: ALL to review the VSA metadata schema in trunk/sql SQL script VSA_MultiframeSchema.sql ACTION: ETWS to check for corresponding essential FITS keys for VSA ingest against those defined for the WSA Continues; awaiting the latest nightly pipeline products from CASU... ACTION: RPB: backup all DB-driven products after UDS DR5. Continues; this can now be done; the team agreed that all science image products (diffs, mosaics, stacks) should be backed up from /disk*/wsa/products/ and RSC asked that /disk05 is included for various bits and pieces. Migration of public DBs to the cluster servers takes priority, so this has had to be delayed (see Hardware below) Actions carried forward from 02/10/09 meeting: ---------------------------------------------- None this week Specific points and new actions: -------------------------------- Project management: PMW and NCH presented the Q4 plan to the team, and after a little discussion it was ratified as sensible. NCH emphasised the need to prioritise VSA over everything now. ACTION: NCH to communicate with some VISTA PS and the SV PIs to finalise some schema details. WFCAM & VISTA updates: UKIDSS ops recommenced on Oct 1st; since then weather not co-operating, but data is trickling in... Comments and issues arising from CASU minutes: WFAU eagerly await new VISTA processed products to shake down things at our end... we hope to get some prior to being suddenly presented with the SV data proper ...! Networking: All quiet on the network interfaces. WSA Operations: All routine ops progressing as usual with the exception of a DB locking problem with the trasnfer utility - ETWS and RPB are investigating. Hardware and Systems: NCH asked that a new interface card be ordered for linux-side backups since the old one was swiped and put in the new cluster system. ACTION: RPB/PMW to spec/order a SCSI card to enable tape from the linux servers Software: RSC reported: "Following the creation of the new release_12 branch of the software to release NJC's schema and software changes to curate the DXS for DR6, I've been checking through the other curation use cases. I realised that the deblended detection parameters fix for the non-surveys will force us to apply his schema changes to the non-survey tables beforehand. This meant I had to implement a new, fourth, mode of operation for ProgrammeBuilder that combines a re-parse of the schema files using the existing set-up together with a database table schema update. This is ready to be applied now before the deblend fix for the non-surveys. I also discovered that this release introduced a silent bug in shallow survey source merging that would have lead to some source table columns having incorrect values. This is now fixed and a warning put in place to avoid such bugs in future." Survey Data Release: LAS, GCS and DXS DR6 all checked and ready to go now. Creation of the DR6 phase-1 DB on the public servers will start today; should finish sometime tomorrow so release announcement early next week. Some fraught communications with GPS Survey Head Phil Lucas concerning presentation of filtered images in the archive user interface image applications have been received over the last week or so. After some discussions, MAR gamely undertook to change the UI so that it does what Phil wants, i.e. essentially hide the filtered images by presenting cut-outs, colour combos etc created from the original unfiltered image despite catalogue attributes coming from the filtered. (Filtered will still be available via suitable settings in the Archive Listing facility). Non-survey Data Release: RPB noted copying of all release DBs over to the new cluster public servers; new releases await the DB deblend fix and associated schema updates. Astrogrid deployment & Data Analysis services: Nothing new to report this week. Miscellaneous: Nothing else this week. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================